# Power sector effects of alternative options for de-fossilizing heavy-duty vehicles: go electric, and charge smartly

### v3

@@ 2.3 Time series reveal differences in flexibility characteristics
Figure 4: Electricity time series sample of five exemplary days (unit: GW): “HDV electricity demand” illustrates electricity flowing to BEV or ERS-BEV batteries, including amounts later charged back via V2G, or electricity needs of hydrogen and PtL supply chains; “HDV discharging to the grid” are V2G power flows from BEV or ERS-BEV to the grid; and “Residual load” is the net load of the electricity sector after considering the feed-in potential of variable renewables. The samples start on a Saturday. For simplification, we assume that both Saturday and Sunday are completely truck-free. BEV Flex and ERS-BEV Flex without V2G are omitted in the graph, as charging patterns are very similar to the shown V2G cases. Additional time series are shown in the SI.
@@
Flexibly operated BEV trucks (BEV Flex V2G) are able to charge their batteries in hours of low residual load, as illustrated in Figure 4. Such periods generally go along with low wholesale market prices (compare Figure SI.9). This enables BEV to utilize renewable surplus energy (i.e., negative residual load) to a substantial extent. They also make use of the V2G option and feed back renewable surplus energy to the grid, whenever the battery capacity is not needed for driving. This is particularly visible during summer during weekends, when trucks are not driving, but assumed to have a grid connection (top right panel in Figure 4). Here, BEV with V2G charge their batteries with solar PV surplus energy on Saturday afternoon, and feed it back to the grid in the night between Saturday and Sunday. This is different in the following days, as HDV are used between Monday and Friday, and much less battery capacity is available for V2G. If BEV charging is not optimized, but follows a pre-determined pattern (BEV Inflex), charging profiles are less peaky and much more balanced. BEV are not able to make particular use of cheap renewable surplus electricity in this case, but carry out a substantial part of their charging in hours with positive residual load.
@@
ERS-BEV follow similar patterns as non-catenary BEV. Yet, their smaller batteries make them temporally less flexible, so they can make less use of renewable surplus energy and have to draw more electricity from the grid during hours with positive residual load. Their potential for feeding electricity back to the grid is thus also much smaller than in the case of pure BEV. ERS-BEV sometimes also charge their batteries during driving (compare Figure SI.10).
@@
Hydrogen and PtL supply chains show very different patterns of electricity use. Centralized hydrogen supply chains (FCEV Centralized), which come with large-scale hydrogen storage capacities, generally have a flat consumption profile in many hours, as their high fixed costs make it optimal to use them with high full load hours. This limits their ability to make use of renewable surplus energy. Yet, they can use the temporal flexibility provided by large-scale hydrogen storage to reduce electricity consumption in hours of high positive residual load, i.e., high prices. Decentralized hydrogen supply chains (FCEV On-Site) have an electricity use profile that largely follows the hydrogen demand profile, driven by limited on-site hydrogen storage capacity. Accordingly, their potential for avoiding electricity consumption in hours of high residual load is lower. The PtL supply chain (ICEV PtL) also comes with large-scale storage and thus has a similar pattern as the centralized hydrogen option. The load, however, is generally higher because of higher overall energy consumption of PtL versus hydrogen. This goes along with larger investments in electrolyzers, which are 24.4 GW in the ICEV PtL scenario, as compared to 13.5 GW for FCEV Centralized (section SI.3.2).
@@ 3.1 Limitations and avenues for future research
First, we deliberately focus on a detailed analysis of the power sector implications of alternative HDV electrification options to address this particular research gap. Unlike other studies, we do neither aim to compare total costs of ownership (nykvist_feasibility_2021; hunter_spatial_2021; samet_2024), nor the overall system costs of different options. Including the latter in our power sector-focused analysis would also require considering infrastructure costs related to catenary lines, chargers, or filling stations, as well as differences in vehicle costs. Such future costs are highly uncertain, so Monte Carlo approaches or other methods to deal with parameter uncertainty would have to be used. Note that the differences in the costs of vehicles and related charging infrastructure might in fact partially outweigh the differences in power sector costs identified here, especially the small differences between pure BEV and ERS-BEV. Yet, the very large power sector cost differences between (ERS-)BEV on the one hand, and FCEV or PtL trucks on the other, would require the latter to be substantially cheaper in terms of vehicle and infrastructure costs in order to break even, which appears unlikely.
@@
While focusing on the power sector, we further assume extensive charging and overhead line infrastructure to be available for BEV and ERS-BEV, both in depots and en route, and that vehicles are always connected to the grid when idling. In reality, the availability of grid connections is likely to be lower for a variety of reasons, which would decrease the flexibility potentials of optimized charging. In sensitivity analyses, we show that decreasing charging availability at depots reduces the potential for flexible charging and particularly for vehicle-to-grid operations of BEV HDV (SI.4.3). This appears to be in line with the finding that the flexibility benefits of very high levels of grid connection may not necessarily outweigh related infrastructure costs (brinkel_2020). In further sensitivities, we conversely find that higher assumptions on fast-charging power ratings do not have a major effect on results.
@@
Another simplification of our numerical analysis is the assumption of perfect foresight, which results in an upper limit of the flexibility benefits that can be realized by optimized (dis-)charging of BEV and ERS-BEV. We also abstract from including electric passenger vehicles or other future sector coupling options, which may provide flexibility to the power sector and compete with electric trucks for electricity use in times of high renewable availability. In particular, power-to-heat technologies may be able to efficiently exploit hours of high renewable generation and low energy prices, as heat storage is relatively cheap compared to electricity storage (bloess_2018). In contrast, we may underestimate temporal flexibility constraints of the PtL production process.
@@
To facilitate numerically efficient solutions and keep results tractable, we further assume electricity generation capacities to be fixed in Germany’s neighboring countries. At the same time, the modeled HDV fleets only cover domestic truck transport in Germany. This may lead to an overestimation of the power sector flexibility supplied by Germany’s neighbors for integrating electric trucks in Germany. Likewise, increasing the spatial resolution within a country could raise additional insights. Taking into account transmission grid bottlenecks and differences in regional potentials for wind and solar power, it may be more beneficial to charge HDV trucks in some regions than in others. We further abstract from comparing different options for the distribution of hydrogen to filling stations, which has been done in detail in previous work (stoeckl_2021).
@@ 3.2 Differences in temporal flexibility and energy efficiency drive effects
[…]
Additional sensitivity analyses show that results are generally robust. For instance, Germany is modeled as an electric island in a sensitivity analysis, i.e., the flexibility benefits of the European interconnection are neglected (section SI.4.1). In such a case, flexible BEV with V2G are even more beneficial, and inflexible BEV operations lead to slightly higher cost increases because an isolated power sector is more flexibility-constrained. Further, results are robust against the assumption that the potential for wind power expansion is limited, and solar PV accordingly plays a larger role (section SI.4.2).
@@ 3.3 Direct electrification and smart charging preferable
From a pure power sector perspective, direct electrification of the truck fleet would thus be clearly preferable. Aside from the cost benefits determined here, this appears to be even more important as the utilization of renewable energy sources is unlikely to grow exponentially (hansen_2017), and renewable growth rates in fact have been smaller than what would be required in $1.5^{\circ}C$-compatible scenarios cherp_2021. A similar argument has been made for the global scale-up of electrolysis (odenweller_2022). All of this calls for energy-efficient, direct electrification options of HDV wherever possible.
Moreover, smart charging of BEV or ERS-BEV is highly desirable, as it provides temporal flexibility which leads to lower power sector costs and CO2 emissions, compared to non-optimized charging. This corroborates the findings of Pickering et al., who also highlight the benefits of controlled vehicle charging in decarbonized and largely renewable future European energy systems pickering_diversity_2022. If feeding back electricity to the grid via V2G is possible, this further lowers the costs and increases the optimal shares of solar PV.
@@ 4.1 The power sector model DIETER
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER). It is a linear program that minimizes power sector costs by optimizing capacity and dispatch decisions for a full year in an hourly resolution zerrahn_long-run_2017; gaete_2021. Its objective function includes fixed an variable costs of all electricity generation and storage technologies, electrolysis and PtL plants, as well as hydrogen or e-fuels transportation. It does not include the costs of charging or catenary infrastructure, hydrogen filling stations, or PtL filling stations. Accordingly, the power sector cost figures provided above do not include the costs of HDV electrification infrastructure. We further do not consider the option of hydrogen imports, as these are likely to be unavailable at scale by 2030. In general, the global scaling up of green hydrogen supply remains uncertain odenweller_2022.
@@
Endogenous model variables include power sector costs, optimal generation and electricity storage capacities (Germany) and their hourly use (all countries), hourly decisions for HDV charging and discharging, as well as the capacity and operational decisions of electrolysis and PtL generation and storage infrastructure. In addition, we interpret the marginals of the hourly energy balance as wholesale prices (brown_decreasing_2021).
@@
Exogenous model inputs include fixed and variable costs of all electricity generation and storage technologies, efficiency parameters, as well as time-series of variable renewable energy availability and electric load. We further include HDV driving profiles as explained in the following section 4.2 and section SI.1. In the case of inflexible HDV charging (BEV Inflex, ERS-BEV Inflex), we assume that the vehicles always start charging as soon as an opportunity arises, and that vehicles batteries are fully charged by the time the next trip starts, if possible. That is, the charging power is lower, the longer a vehicle is connected to the grid. This resembles the “balanced” charging profile defined in gaete_2021_emobpy.
@@
The geographic scope of the model version used here includes Germany and its neighboring countries plus Italy. In order to reduce numerical complexity and improve tractability, we allow for endogenous generation capacity investment only in Germany, and fix the power plant portfolio for the other countries to values derived from ENTSO-E’s Ten Year Network Development Plans tyndp_2018; tyndp_2020. We further assume upper limits for investments into fossil generation technologies according to the main scenario for 2030 from the German Federal Grid Development plan (NEP, nep_2018).
@@
Different versions of the model have been used in various earlier studies to investigate various aspects of renewable energy integration, electricity storage, and sector coupling (e.g.,  schill_2018; schill_joule_2020; stoeckl_2021; gils_2022). The DIETER model is available open-source. [^7] General DIETER repo: https://gitlab.com/diw-evu/dieter_public. The particular model version used for the present study, together with the complete input data set (including BEV and ERS-BEV profiles), is available under permissive licenses in a dedicated repository. This also includes model results in Jupyter notebook and html formats: https://gitlab.com/diw-evu/projects/my_eroads. A general model documentation is available here: https://diw-evu.gitlab.io/dieter_public/dieterpy/.
@@ 4.2 Traffic data of heavy-duty vehicles
We generate synthetic truck usage patterns that are intended to approximate the German fleet of HDV larger than 26 tons. The main data source for the usage patterns is the traffic model PTV Validate (ptv_2023), from which we extract a database of daily truck trips in domestic German road freight transport. In order to derive typical daily driving profiles, the following steps are performed (for more detailed information, see SI):
@@
For FCEV and PtL trucks, only aggregate fuel demand time series are required. These are based on the same empirical traffic data from German highways (bast_2023). For FCEV, we assume a specific hydrogen demand of 6.8 kg/100km, and for ICEV-PtL a specific Diesel demand of 27.1 l/100km.
@@ Acknowledgments
Resource availability
The DIETER model version used for the present study, including input data as well as model results, are available under permissive licenses in a public repository: https://gitlab.com/diw-evu/projects/my_eroads.

### v4

@@ 2.3 Time series reveal differences in flexibility characteristics
Figure 4: Electricity time series sample of five exemplary days (unit: GW): “HDV electricity demand” illustrates electricity flowing to BEV or ERS-BEV batteries, including amounts later charged back via V2G, or electricity needs of hydrogen and PtL supply chains; “HDV discharging to the grid” are V2G power flows from BEV or ERS-BEV to the grid; and “Residual load” is the net load of the electricity sector after considering the feed-in potential of variable renewables. The samples start on a Saturday. For simplification, we assume that both Saturday and Sunday are completely truck-free. BEV Flex and ERS-BEV Flex without V2G are omitted in the graph, as charging patterns are very similar to the shown V2G cases. Additional time series are shown in the SI, Figures S.10S.12S.13.
@@
Flexibly operated BEV trucks (BEV Flex V2G) are able to charge their batteries in hours of low residual load, as illustrated in Figure 4. Such periods generally go along with low wholesale market prices (compare Figure S.11). This enables BEV to utilize renewable surplus energy (i.e., negative residual load) to a substantial extent. They also make use of the V2G option and feed back renewable surplus energy to the grid, whenever the battery capacity is not needed for driving. This is particularly visible during summer during weekends, when trucks are not driving, but assumed to have a grid connection (top right panel in Figure 4). Here, BEV with V2G charge their batteries with solar PV surplus energy on Saturday afternoon, and feed it back to the grid in the night between Saturday and Sunday. This is different in the following days, as HDV are used between Monday and Friday, and much less battery capacity is available for V2G. If BEV charging is not optimized, but follows a pre-determined pattern (BEV Inflex), charging profiles are less peaky and much more balanced. BEV are not able to make particular use of cheap renewable surplus electricity in this case, but carry out a substantial part of their charging in hours with positive residual load.
@@
ERS-BEV follow similar patterns as non-catenary BEV. Yet, their smaller batteries make them temporally less flexible, so they can make less use of renewable surplus energy and have to draw more electricity from the grid during hours with positive residual load. Their potential for feeding electricity back to the grid is thus also much smaller than in the case of pure BEV. ERS-BEV sometimes also charge their batteries during driving (compare Figure S.12).
@@
Hydrogen and PtL supply chains show very different patterns of electricity use. Centralized hydrogen supply chains (FCEV Centralized), which come with large-scale hydrogen storage capacities, generally have a flat consumption profile in many hours, as their high fixed costs make it optimal to use them with high full load hours. This limits their ability to make use of renewable surplus energy. Yet, they can use the temporal flexibility provided by large-scale hydrogen storage to reduce electricity consumption in hours of high positive residual load, i.e., high prices. Decentralized hydrogen supply chains (FCEV On-Site) have an electricity use profile that largely follows the hydrogen demand profile, driven by limited on-site hydrogen storage capacity. Accordingly, their potential for avoiding electricity consumption in hours of high residual load is lower. The PtL supply chain (ICEV PtL) also comes with large-scale storage and thus has a similar pattern as the centralized hydrogen option. The load, however, is generally higher because of higher overall energy consumption of PtL versus hydrogen. This goes along with larger investments in electrolyzers, which are 24.4 GW in the ICEV PtL scenario, as compared to 13.5 GW for FCEV Centralized (section S.2.2).
@@ 3.1 Limitations and avenues for future research
First, we deliberately focus on a detailed analysis of the power sector implications of alternative HDV electrification options to address this particular research gap. Unlike other studies, we do neither aim to compare total costs of ownership ([nykvist_feasibility_2021, hunter_spatial_2021, samet_2024]), nor the overall system costs of different options. Including the latter in our power sector-focused analysis would also require considering infrastructure costs related to catenary lines, chargers, or filling stations, as well as differences in vehicle costs. Such future costs are highly uncertain, so Monte Carlo approaches or other methods to deal with parameter uncertainty would have to be used. Note that the differences in the costs of vehicles and related charging infrastructure might in fact partially outweigh the differences in power sector costs identified here, especially the small differences between pure BEV and ERS-BEV. Yet, the very large power sector cost differences between (ERS-)BEV on the one hand, and FCEV or PtL trucks on the other, would require the latter to be substantially cheaper in terms of vehicle and infrastructure costs in order to break even, which appears unlikely.
@@
While focusing on the power sector, we further assume extensive charging and overhead line infrastructure to be available for BEV and ERS-BEV, both in depots and en route, and that vehicles are always connected to the grid when idling. In reality, the availability of grid connections is likely to be lower for a variety of reasons, which would decrease the flexibility potentials of optimized charging. In sensitivity analyses, we show that decreasing charging availability at depots reduces the potential for flexible charging and particularly for vehicle-to-grid operations of BEV HDV (S.2.8). This appears to be in line with the finding that the flexibility benefits of very high levels of grid connection may not necessarily outweigh related infrastructure costs [brinkel_2020]. In further sensitivities, we conversely find that higher assumptions on fast-charging power ratings do not have a major effect on results (Figures S.20S.21S.22).
@@
Another simplification of our numerical analysis is the assumption of perfect foresight, which results in an upper limit of the flexibility benefits that can be realized by optimized (dis-)charging of BEV and ERS-BEV. We also abstract from including electric passenger vehicles or other future sector coupling options, which may provide flexibility to the power sector and compete with electric trucks for electricity use in times of high renewable availability. In particular, power-to-heat technologies may be able to efficiently exploit hours of high renewable generation and low energy prices, as heat storage is relatively cheap compared to electricity storage [bloess_2018]. In contrast, we may underestimate temporal flexibility constraints of the PtL production process.
@@
To facilitate numerically efficient solutions and keep results tractable, we further assume electricity generation capacities to be fixed in Germany’s neighboring countries. At the same time, the modeled HDV fleets only cover domestic truck transport in Germany. This may lead to an overestimation of the power sector flexibility supplied by Germany’s neighbors for integrating electric trucks in Germany. Likewise, increasing the spatial resolution within a country could raise additional insights. Taking into account transmission grid bottlenecks and differences in regional potentials for wind and solar power, it may be more beneficial to charge HDV trucks in some regions than in others. We further abstract from comparing different options for the distribution of hydrogen to filling stations, which has been done in detail in previous work [stoeckl_2021].
@@ 3.2 Differences in temporal flexibility and energy efficiency drive effects
Alternative options of (in-)directly electrifying HDV fleets further entail different optimal generation capacity mixes. The additional electricity for battery-electric HDV is largely supplied by solar PV in the lowest-cost solution, particularly if V2G is also available. In contrast, centralized hydrogen and PtL supply chains go along with higher optimal shares of wind power, and also lead to a higher use of non-renewable electricity generators. The latter translates into higher direct carbon emissions of these technologies. However, emission outcomes depend on the assumed boundary conditions, such as capacity bounds for non-renewable generators and CO2 prices.
Additional sensitivity analyses show that results are generally robust. For instance, Germany is modeled as an electric island in a sensitivity analysis, i.e., the flexibility benefits of the European interconnection are neglected (section S.2.6, Figures S.14S.15S.16). In such a case, flexible BEV with V2G are even more beneficial, and inflexible BEV operations lead to slightly higher cost increases because an isolated power sector is more flexibility-constrained. Further, results are robust against the assumption that the potential for wind power expansion is limited, and solar PV accordingly plays a larger role (section S.2.7, Figures S.17S.18S.19).
@@ 3.3 Direct electrification and smart charging preferable
From a pure power sector perspective, direct electrification of the truck fleet would thus be clearly preferable. Aside from the cost benefits determined here, this appears to be even more important as the utilization of renewable energy sources is unlikely to grow exponentially [hansen_2017], and renewable growth rates in fact have been smaller than what would be required in $1.5^{\circ}C$-compatible scenarios [cherp_2021]. A similar argument has been made for the global scale-up of electrolysis [odenweller_2022]. All of this calls for energy-efficient, direct electrification options of HDV wherever possible.
Moreover, smart charging of BEV or ERS-BEV is highly desirable, as it provides temporal flexibility which leads to lower power sector costs and CO2 emissions, compared to non-optimized charging. This corroborates the findings of Pickering et al., who also highlight the benefits of controlled vehicle charging in decarbonized and largely renewable future European energy systems [pickering_diversity_2022]. If feeding back electricity to the grid via V2G is possible, this further lowers the costs and increases the optimal shares of solar PV.
@@ 4 Experimental procedures
4.1 Resource availability
4.1.1 Lead contact
Further information and requests for resources should be directed to the lead contact, Wolf-Peter Schill (wschill@diw.de)
4.1.2 Materials availability
This study did not generate new unique reagents.
4.1.3 Data and code availability
All data and code are available under permissive licenses in public repositories. The general repository of the DIETER model can be found here: https://gitlab.com/diw-evu/dieter_public. The particular model version used for the present study, together with the complete input data set (including BEV and ERS-BEV profiles), is available under permissive licenses in a dedicated repository. This also includes model results in Jupyter notebook and html formats: https://gitlab.com/diw-evu/projects/my_eroads. A general model documentation is available here: https://diw-evu.gitlab.io/dieter_public/dieterpy/.
@@ 4.2 The power sector model DIETER
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER). It is a linear program that minimizes power sector costs by optimizing capacity and dispatch decisions for a full year in an hourly resolution [zerrahn_long-run_2017, gaete_2021]. Its objective function includes fixed an variable costs of all electricity generation and storage technologies, electrolysis and PtL plants, as well as hydrogen or e-fuels transportation. It does not include the costs of charging or catenary infrastructure, hydrogen filling stations, or PtL filling stations. Accordingly, the power sector cost figures provided above do not include the costs of HDV electrification infrastructure. We further do not consider the option of hydrogen imports, as these are likely to be unavailable at scale by 2030. In general, the global scaling up of green hydrogen supply remains uncertain [odenweller_2022].
@@
Endogenous model variables include power sector costs, optimal generation and electricity storage capacities (Germany) and their hourly use (all countries), hourly decisions for HDV charging and discharging, as well as the capacity and operational decisions of electrolysis and PtL generation and storage infrastructure. In addition, we interpret the marginals of the hourly energy balance as wholesale prices [brown_decreasing_2021].
@@
Exogenous model inputs include fixed and variable costs of all electricity generation and storage technologies, efficiency parameters, as well as time-series of variable renewable energy availability and electric load. We further include HDV driving profiles as explained in the following section 4.3 and section S.1.1. In the case of inflexible HDV charging (BEV Inflex, ERS-BEV Inflex), we assume that the vehicles always start charging as soon as an opportunity arises, and that vehicles batteries are fully charged by the time the next trip starts, if possible. That is, the charging power is lower, the longer a vehicle is connected to the grid. This resembles the “balanced” charging profile defined in [gaete_2021_emobpy].
@@
The geographic scope of the model version used here includes Germany and its neighboring countries plus Italy. In order to reduce numerical complexity and improve tractability, we allow for endogenous generation capacity investment only in Germany, and fix the power plant portfolio for the other countries to values derived from ENTSO-E’s Ten Year Network Development Plans [tyndp_2018, tyndp_2020]. We further assume upper limits for investments into fossil generation technologies according to the main scenario for 2030 from the German Federal Grid Development plan (NEP, [nep_2018]).
@@
Different versions of the model have been used in various earlier studies to investigate various aspects of renewable energy integration, electricity storage, and sector coupling [e.g.,  schill_2018, schill_joule_2020, stoeckl_2021, gils_2022]. The DIETER model is available open-source.
@@ 4.3 Traffic data of heavy-duty vehicles
We generate synthetic truck usage patterns that are intended to approximate the German fleet of HDV larger than 26 tons. The main data source for the usage patterns is the traffic model PTV Validate [ptv_2023], from which we extract a database of daily truck trips in domestic German road freight transport. In order to derive typical daily driving profiles, the following steps are performed (for more detailed information, see SI):
@@
For FCEV and PtL trucks, only aggregate fuel demand time series are required. These are based on the same empirical traffic data from German highways [bast_2023]. For FCEV, we assume a specific hydrogen demand of 6.8 kg/100km, and for ICEV-PtL a specific Diesel demand of 27.1 l/100km.
@@ Acknowledgments
Declaration of interests
The authors declare no competing interests.